OpenStudy (brainzonly):

Car A began a journey from a point at 9 am, traveling at 30 mph. At 10 am car B started traveling from the same point at 40 mph in the same direction as car A. At what time will car B pass car A?

OpenStudy (brainzonly):

@Kevin Got 1 for ya

OpenStudy (brainzonly):

This is what I have so far: Let after t hours the distances D1 traveled by car A => D1 = 30 t Car B starts at 10 am and will therefore have spent one hour less than car A when it passes it. After (t - 1) hours, distance D2 traveled by car B => D2 = 40 (t - 1)

OpenStudy (kevin):

wait

OpenStudy (kevin):

let me solve this

OpenStudy (brainzonly):

gotta go

OpenStudy (kevin):

How many the distance between A and B start position?

OpenStudy (kevin):

Owh... I get it

OpenStudy (kevin):

30.t = 40(t - 1) t = 4 hours So it will be 9 + 4 = 1 PM

OpenStudy (brainzonly):

I'm back. I was on the phone :D I multitasked and found out this: Let after t hours the distances D1 traveled by car A => D1 = 30 t Car B starts at 10 am and will therefore have spent one hour less than car A when it passes it. After (t - 1) hours, distance D2 traveled by car B => D2 = 40 (t - 1) When car B passes car A, they are at the same distance from the starting point and therefore D1 = D2 => 30 t = 40 (t - 1) Solve the equation for t, => 30 t = 40t - 40 => 10 t = 40 => t = 4 => Car B passes car A at = 9 + 4 = 13 pm.

OpenStudy (kevin):

That's a very good question. I need more question like this :D

OpenStudy (brainzonly):

OK! opening new q in a minute